# Privacy Loss of Noise Perturbation via Concentration Analysis of A Product Measure

### v3

@@ 4.2. A New PLRV
Figure 1. Geometric interpretation of noise perturbation.:
@@ 4.5.2. Simulation
(a) $\log_{10}(\delta)$ v.s. $M$ for various tuning parameters k:
(b) $\mathbb{E}[\|\mathbf{n}\|_{2}^{2}]$ v.s. $M$ for various noises:
@@ 5.4. Gradient Perturbation for Non-Convex ERM
Given a gradient norm clipping parameter $C$, the product noise-based DPSGD achieves ($\epsilon,\delta$)-DP for each gradient descend step, if ${\sigma_{M}}=\frac{2\sqrt{2}C}{k}^{\frac{2}{M}}\left(\frac{M}{4}+\frac{3}{2}\right)^{\left(\frac{1}{2}+\frac{2}{M}\right)}{\epsilon e^{\left(\frac{1}{2}+\frac{1}{M}\right)}}$. Note $\delta$ is determined by $k$ and the dimension of the problem $M$ (ref. (13)).Given a data sampling probability $p$ and training steps $T$, the cumulative privacy loss of our product noise-based DPSGD is evaluated using privacy amplification via subsampling (see Theorem 4 in Section 3) followed by a distribution-independent composition (he2021tighter) (see Theorem 5 in Section 3).
@@ 6.1. Case Study I: Output Perturbation
(a) MNIST :
(b) RCV1 :
@@ 6.2. Case Study II: Objective Perturbation
(a) MNIST :
(b) Synthetic-H :
(c) Real-sim:
(d) RCV1 :
@@
(a) Test Accuracy vs Epochs:
(b) Training Accuracy vs Epochs:
(c) Epsilon vs Epochs:
@@
(a) Test Accuracy vs Epochs:
(b) Training Accuracy vs Epochs:
(c) Epsilon vs Epochs:
@@ A.6. Solving for $\delta=\min\limits_{q}\delta_{1}\delta_{2}$
(a) $\delta$ versus $q$ ($M=10^{6}$):
(b) $\delta$ versus $q$ ($M=10^{8}$):
@@ E.2. Supplemental Experiments for Output Perturbation
(a) Adult :
(b) KDDCup99 :
(c) Synthetic-H :
(d) Real-sim :
@@
Table 6. $\ell_{2}$ error error of Output Perturbation on LR.
@@
Table 7. FPR error of Output Perturbation on LR.
@@
(a) Adult:
(b) KDDCup99 :
(c) MNIST :
(d) Synthetic-H :
(e) Real-sim :
(f) RCV1 :
@@
(a) Adult :
(b) KDDCup99 :
(c) MNIST:
(d) Synthetic-H :
(e) Real-sim:
(f) RCV1 :
@@
(a) Adult :
(b) KDDCup99 :
(c) MNIST:
(d) Synthetic-H :
(e) Real-sim :
(f) RCV1 :
@@ E.3. Supplemental Experiments for Objective Perturbation on Huber SVM
(a) MNIST :
(b) Synthetic-H :
(c) Real-sim:
(d) RCV1 :
@@ E.5. Gradient Perturbation on Other Datasets
(a) Test Accuracy vs Epochs:
(b) Training Accuracy vs Epochs:
(c) Epsilon vs Epochs:
@@
(a) Test Accuracy vs Epochs:
(b) Training accuracy vs Epochs:
(c) Epsilon vs Epochs:
@@
(a) Test Accuracy vs Epochs:
(b) Training Accuracy vs Epochs:
(c) Epsilon vs Epochs:

### v4

@@ 4.2. A New PLRV
Refer to caption Geometric interpretation of differentially private output perturbation.
Figure 1. Geometric interpretation of noise perturbation.
@@ 4.5.2. Simulation
plot comparing product Gaussian noise and classical Gaussian noise as dimension increases.
@@
Histogram showing the squared magnitude of noise for the proposed noise mechanism.
@@
Refer to caption Plot of log10(delta) versus M.
(a) $\log_{10}(\delta)$ v.s. $M$ for various tuning parameters k
Refer to caption Plot of log10(expected squared noise magnitude) versus dimension M for various noises.
(b) $\mathbb{E}[\|\mathbf{n}\|_{2}^{2}]$ v.s. $M$ for various noises
@@ 5.4. Gradient Perturbation for Non-Convex ERM
Given a data sampling probability $p$ and training steps $T$, the cumulative privacy loss of our product noise-based DPSGD is evaluated using privacy amplification via subsampling (see Theorem 4 in Appendix A.2) followed by a distribution-independent composition (he2021tighter) (see Theorem 5 in Appendix A.2).
@@ 6.1. Case Study I: Output Perturbation
Refer to caption Accuracy results of output perturbation for MNIST dataset (Low-Dim).
(a) MNIST
Refer to caption Accuracy results of output perturbation for RCV1 dataset (High-Dim).
(b) RCV1
@@ 6.2. Case Study II: Objective Perturbation
Refer to caption Graph showing accuracy versus epsilon for MNIST dataset (Low-Dim) under objective perturbation for LR.
(a) MNIST
Refer to caption Graph showing accuracy versus epsilon for Synthetic-H dataset (High-Dim) under objective perturbation for LR.
(b) Synthetic-H
Refer to caption Graph showing accuracy versus epsilon for Real-sim dataset (High-Dim) under objective perturbation for LR.
(c) Real-sim
Refer to caption Graph showing accuracy versus epsilon for RCV1 dataset (High-Dim) under objective perturbation for LR.
(d) RCV1
@@
Refer to caption Graph showing test accuracy results on the MNIST dataset using DPSGD.
(a) Test Accuracy vs Epochs
Refer to caption Graph showing Training accuracy results on the MNIST dataset using DPSGD.
(b) Training Accuracy vs Epochs
Refer to caption Graph comparing epsilon values on the MNIST dataset under DPSGD.
(c) Epsilon vs Epochs
@@
Refer to caption Graph showing test accuracy results on the CIFAR-10 dataset using DPSGD.
(a) Test Accuracy vs Epochs
(b) Training Accuracy vs Epochs
Training accuracy curve of the CIFAR-10 dataset with DPSGD.
Refer to caption Graph comparing epsilon values on the CIFAR-10 dataset under DPSGD.
(c) Epsilon vs Epochs
@@ A.6. Solving for $\delta=\min\limits_{q}\delta_{1}\delta_{2}$
Refer to caption Graph showing delta versus q for M=10^6.
(a) $\delta$ versus $q$ ($M=10^{6}$)
Refer to caption Graph showing delta versus q for M=10^8.
(b) $\delta$ versus $q$ ($M=10^{8}$)
@@ E.2. Supplemental Experiments for Output Perturbation
Refer to caption Accuracy results of output perturbation for Adult dataset.
(a) Adult
Refer to caption Accuracy results of output perturbation for KDDCup99 dataset.
(b) KDDCup99
Refer to caption Accuracy results of output perturbation for Synthetic-H dataset.
(c) Synthetic-H
Refer to caption Accuracy results of output perturbation for Real-sim dataset .
(d) Real-sim
@@
Table 6. $\ell_{2}$ error of Output Perturbation on LR.
@@
Table 7. FPR of Output Perturbation on LR.
@@
Refer to caption Accuracy results of output perturbation for Adult dataset .
(a) Adult
Refer to caption Accuracy results of output perturbation for KDDCup99 dataset .
(b) KDDCup99
Refer to caption Accuracy results of output perturbation for MNIST dataset .
(c) MNIST
Refer to caption Accuracy results of output perturbation for Synthetic-H dataset .
(d) Synthetic-H
Refer to caption Accuracy results of output perturbation for Real-sim dataset .
(e) Real-sim
Refer to caption Accuracy results of output perturbation for RCV1 dataset .
(f) RCV1
@@
Refer to caption Accuracy results of Private Strongly Convex Permutation-based for Adult dataset .
(a) Adult
Refer to caption Accuracy results of Private Strongly Convex Permutation-based for KDDCup99 dataset .
(b) KDDCup99
Refer to caption Accuracy results of Private Strongly Convex Permutation-based for MNIST dataset (Low-Dim).
(c) MNIST
Refer to caption Accuracy results of o Private Strongly Convex Permutation-based for Synthetic-H dataset .
(d) Synthetic-H
Refer to caption Accuracy results of Private Strongly Convex Permutation-based for Real-sim dataset.
(e) Real-sim
Refer to caption Accuracy results of Private Strongly Convex Permutation-based for RCV1 dataset (High-Dim).
(f) RCV1
@@
Refer to caption Accuracy results of output perturbation for Adult dataset .
(a) Adult
Refer to caption Accuracy results of output perturbation for KDDCup99 dataset .
(b) KDDCup99
Refer to caption Accuracy results of output perturbation for MNIST dataset.
(c) MNIST
Refer to caption Accuracy results of output perturbation for Synthetic-H dataset .
(d) Synthetic-H
Refer to caption Accuracy results of output perturbation for Real-sim dataset .
(e) Real-sim
Refer to caption Accuracy results of output perturbation for RCV1 dataset.
(f) RCV1
@@ E.3. Supplemental Experiments for Objective Perturbation on Huber SVM
Refer to caption Graph showing accuracy versus epsilon for MNIST dataset under objective perturbation for Huber SVM.
(a) MNIST
Refer to caption Graph showing accuracy versus epsilon for Synthetic-H dataset under objective perturbation for Huber SVM.
(b) Synthetic-H
Refer to caption Graph showing accuracy versus epsilon for Real-sim dataset under objective perturbation for Huber SVM.
(c) Real-sim
Refer to caption Graph showing accuracy versus epsilon for RCV1 dataset under objective perturbation for Huber SVM.
(d) RCV1
@@ E.5. Gradient Perturbation on Other Datasets
Refer to caption Graph showing test accuracy results on the Adult dataset using DPSGD.
(a) Test Accuracy vs Epochs
(b) Training Accuracy vs Epochs
Training accuracy curve of the Adult dataset with DPSGD.
Refer to caption Graph comparing epsilon values on the Adult dataset under DPSGD.
(c) Epsilon vs Epochs
@@
Refer to caption Graph showing test accuracy results on the IMDb dataset using DPSGD.
(a) Test Accuracy vs Epochs
(b) Training accuracy vs Epochs
Training accuracy curve of the IMDb dataset with DPSGD.
Refer to caption Graph comparing epsilon values on the IMDb dataset under DPSGD.
(c) Epsilon vs Epochs
@@
Refer to caption Graph showing test accuracy results on the MovieLens dataset using DPSGD.
(a) Test Accuracy vs Epochs
(b) Training Accuracy vs Epochs
Training accuracy curve of the MovieLens dataset with DPSGD.
Refer to caption Graph comparing epsilon values on the MovieLens dataset under DPSGD.
(c) Epsilon vs Epochs